# Adaptive Bounding Box Uncertainties via Two-Step Conformal Prediction

### v1

@@ Abstract
Quantifying a model’s predictive uncertainty is essential for safety-critical applications such as autonomous driving. We consider quantifying such uncertainty for multi-object detection. In particular, we leverage conformal prediction to obtain uncertainty intervals with guaranteed coverage for object bounding boxes. One challenge in doing so is that bounding box predictions are conditioned on the object’s class label. Thus, we develop a novel two-step conformal approach that propagates uncertainty in predicted class labels into the uncertainty intervals for the bounding boxes. This broadens the validity of our conformal coverage guarantees to include incorrectly classified objects, ensuring their usefulness when maximal safety assurances are required. Moreover, we investigate novel ensemble and quantile regression formulations to ensure the bounding box intervals are adaptive to object size, leading to a more balanced coverage across sizes. Validating our two-step approach on real-world datasets for 2D bounding box localization, we find that desired coverage levels are satisfied with actionably tight predictive uncertainty intervals.
@@ 1 Introduction
Safety-critical applications in domains such as autonomous transportation [36, 66] and mobile robotics [35] benefit greatly from accurate estimates of the model’s predictive uncertainty. Yet one obstacle to principled uncertainty quantification (UQ) for computer vision is the pervasive use of deep neural networks – which are often unamenable to traditional techniques for UQ. The framework of Conformal Prediction (CP) [62, 52, 2] enables a form of distribution-free UQ that is agnostic to the predictive model’s structure, rendering it well-suited for such ‘black-box’ models.
In this work, we propose a CP framework designed to quantify predictive uncertainties in multi-object detection tasks with multiple classes (see Figure 2). CP allows us to produce computationally cheap, post-hoc distribution-free prediction intervals, which come equipped with a coverage guarantee for the true bounding boxes of previously unseen objects (of known classes). Specifically, we provide users with the following safety assurance: “The conformal prediction interval covers the object’s true bounding box with probability $(1-\alpha)$ for any known object class", where $\alpha$ is an acceptable margin of error. Such a guarantee can, in the context of autonomous driving, help certify collision avoidance by steering clear of the outer-interval limits (see Figure 1), or in the case of robot picking, enforce cautious handling by demarcating a reliable grasping zone via the inner-interval limits.
@@
Employing strategies based on ensembling and quantile regression, we ensure that the obtained intervals are adaptive to object size: they may grow or shrink in individual dimensions to account for object variability and prediction difficulty. A challenge to the desired assurance is that constructed intervals rely on the model’s predicted class labels, which may be potentially erroneous. We thus introduce an additional conformal step over the class labels, shielding against misclassification and ensuring that downstream coverage is satisfied. That is, our two-step conformal pipeline remains theoretically and empirically valid regardless of the underlying object detector’s predictive performance for either class labels or box coordinates – the incurred costs are solely reflected in the obtained prediction interval sizes. In the experiments, we apply our methodology to multiple classes on several real-world 2D object detection datasets. We obtain bounding box prediction intervals that adhere to the desired guarantee, and are both adaptive and practically useful for downstream decision-making.
Figure 2: A diagram of our proposed two-step conformal approach. We compute conformal quantiles for both class labels and box coordinates on calibration data following the CP framework. These are then used on the predictions of a ‘black-box’ object detector for an unseen test sample to (1) form a conformal label set with guarantee (✓) which informs our box quantile choice; and (2) form a conformal prediction interval for the bounding box with guarantee (✓), thus producing a predictive uncertainty estimate with safety assurances.
To summarize, our core contribution is an end-to-end framework for safe bounding box uncertainties which is post-hoc, efficient, and generalizable. In that process, we introduce several original concepts such as (i) ensemble and quantile CP adaptations for object detection, (ii) leveraging strong class-conditional guarantees for multi-class settings, and (iii) proposing a sequential two-step approach that propagates classification uncertainties forward. Such a two-step conformal framework is, to the best of our knowledge, entirely novel for object localization.
@@ 2 Background
We begin by providing background on conformal prediction and the desired coverage guarantee, formalizing our object detection setting, and relating the conformal prediction framework to it.
@@ 2.1 Conformal prediction
Let us consider the most common setting of split CP [42], where we perform a single split to obtain separate calibration data $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}\sim P_{XY}$, as opposed to alternative partitioning schemes [61, 6]. If the general conformal procedure outlined in Algorithm 1 is followed, a coverage guarantee for an unseen test sample $(X_{n+1},Y_{n+1})\sim P_{XY}$ is provided in terms of a prediction set $\hat{C}(X_{n+1})$, where a finite-sample, distribution-free guarantee is given over the event of $\hat{C}(X_{n+1})$ containing $Y_{n+1}$.
Assuming the samples $\mathcal{D}_{cal}\cup\{(X_{n+1},Y_{n+1})\}$ are exchangeable – a relaxation of the i.i.d. assumption – we obtain a probabilistic guarantee that
@@
for some tolerated miscoverage rate $\alpha\in(0,1)$ [52]. The provided guarantee is marginally valid, since it holds on average across any sample $(X_{n+1},Y_{n+1})$ and set $\mathcal{D}_{cal}$ for some fixed distribution $P_{XY}$ over $\mathcal{X}\times\mathcal{Y}$. This is in contrast to the ideal scenario of conditionally valid coverage per input sample $X_{n+1}$, which has been shown to be impossible to achieve in a distribution-free manner [60, 18]. However, recent work on in-between notions of conditionality such as group- [47, 22] and feature-conditional [51] strive towards more granular guarantees.
@@
where $\mathcal{Y}=\{1,\dots,K\}$ are distinct class labels. The class-conditional guarantee in Equation 2 is stronger and implies Equation 1, in that we aim to control some miscoverage level $\alpha$ for samples within each class. In addition, it permits setting individual miscoverage levels $\{\alpha_{y}\}_{y\in\mathcal{Y}}$ per class, if desired, and is robust to changes and imbalances in class proportions [46, 59]. Such a class-conditional guarantee is precisely what we aim to provide.
Classification and regression. Applying CP to a classification task yields conformal label prediction sets $\hat{C}_{L}(X_{n+1})\subseteq\{1,\dots,K\}$ as finite subsets of the $K$ class labels, at a target miscoverage level $\alpha_{L}$. For regression, the sets $\hat{C}_{B}(X_{n+1})\subseteq\mathbb{R}$ take the form of prediction intervals (PIs) on the target domain, at target miscoverage level $\alpha_{B}$.
@@ 2.2 Object detection
We next formalize our multi-object detection setting. Consider an input image $X\in\mathbb{R}^{H\times W\times D}$, where $H$, $W$ and $D$ correspond to image height, width and depth. For each image in $D_{cal}$ we also receive a set of tuples $(c^{1},c^{2},c^{3},c^{4},l)$, where $(c^{1},c^{2},c^{3},c^{4})\in\mathbb{R}^{4}$ are the coordinates indicating an object’s bounding box location in the image, and $l\in\{1,\dots,K\}$ represents the object’s class label.
Each tuple parameterizes an object, with a total of $O(X)$ true objects located in the image. For image $X$ we thus have responses $\{(c^{1},c^{2},c^{3},c^{4},l)_{j}\}_{j=1}^{O(X)}$. Note that the model predicts $\hat{O}(X)$ objects, and it is possible that $O(X)\neq\hat{O}(X)$. We model every object as an individual sample for our CP procedures, i.e., the same input image $X$ will produce multiple calibration samples of shape $(X,(c^{1},c^{2},c^{3},c^{4},l)_{j})$, where $j=1,\dots,O(X)$ denote the object responses.
Object detection model. For our object detector $\hat{f}$, we define two separate output heads. The probabilistic classification head is defined as the map $\hat{f}_{L}:X\mapsto(\hat{\pi}_{1},\dots,\hat{\pi}_{K})$, where $\hat{\pi}_{y}$ is the model’s estimate of the true class probability $\pi_{y}$ of some object in image $X$ belonging to class $y$. The object’s class label is then $l=\arg\max_{y\in\{1,\dots,K\}}\,\hat{\pi}_{y}$. The bounding box regression head, denoted as $\hat{f}_{B}:X\mapsto(c^{1},c^{2},c^{3},c^{4})$, maps to an object’s real-valued bounding box coordinates.
@@ 2.3 Conformal prediction for object detection
Given our multi-object detection setting, we consider a class-conditional CP approach to be particularly meaningful. It is sensible to only leverage information on detected objects of the same class, e.g., class ‘car’, to construct PIs for new objects of that class. In contrast, a general marginal approach will unintuitively also use information from unrelated classes, such as ‘person’ or ‘bicycle’.
We apply CP to the bounding boxes on a per-coordinate basis, previously denoted $(c^{1},c^{2},c^{3},c^{4})$. However, from now on let us consider the generalization to an arbitrary amount of coordinates $c^{k},\,k=1,\dots,m$. If we consider the class label $l$ within each group of objects belonging to a common class as fixed (since all objects share the same label), the response of an individual sample $(X_{i},Y_{i})$ can be interpreted as a realization of the $m$ coordinates only, i.e., we define $Y_{i}:=(c_{i}^{1},\dots,c_{i}^{m})\in\mathbb{R}^{m}$. The desired guarantee from Equation 2 is then re-interpreted as
$\mathbb{P}(\bigcap_{k=1}^{m}(c_{n+1}^{k}\in\hat{C}_{B}^{k}(X_{n+1}))|l_{n+1}=y)\geq 1-\alpha_{B}\quad\forall y\in\mathcal{Y},$ (3)
where components are indexed accordingly for a specific coordinate dimension. For instance, $\hat{C}_{B}^{k}(X_{n+1})$ is the $k$-th coordinate’s prediction interval of an object of class $y$ (its class label $l_{n+1}$ matches $y$) located in image $X_{n+1}$. Applying CP to each coordinate separately gives rise to multiple testing issues, which we address in subsection 4.1.
Practical considerations. Naively applying class-conditional CP to the box coordinates necessitates a correct class label prediction in order to satisfy validity. That is, for Equation 3 to hold, a valid PI construction requires $\hat{l}_{n+1}=l_{n+1}$ for any considered class $y\in\mathcal{Y}$. We alleviate this practically limiting dependence on the model’s classification ability using a conformal set-based classifier in section 5 (see also Figure 2). However, we of course still rely on the model’s general detection abilities – the provided guarantees only hold for true objects that are actually detected (true positives) and do not account for missed objects (false negatives), as also noted by [14, 1].
@@ 3 Related work
Many existing approaches for quantifying uncertainty in bounding box regression leverage standard UQ techniques such as Bayesian inference [12, 24, 68] and loss attenuation [21, 28, 29], or popular approximate Bayesian methods like Monte Carlo Dropout [44, 20, 38, 71] and Deep Ensembles [15, 65, 39]. These can require substantial modifications to the model architecture or training procedure, and do not provide a guarantee or statement of assurance about provided estimation quality. See Feng et al. [16] for a recent survey. A complementary line of work investigates the calibration of object detectors [45, 26, 41], which can empirically benefit our approach by improving the underlying ‘black-box’ model.
Conformal approaches have recently gained traction for computer vision and related tasks, with applications such as image classification [49, 5], geometric pose estimation [70], or tracking and trajectory planning [56, 33, 34, 40]. Yet, the domain remains comparatively unexplored given current surveys [2, 17]. Specific attempts at principled UQ for bounding boxes include using the Probably Approximatly Correct (PAC) framework to produce guarantees by composition of PAC prediction sets at multiple modelling stages [31, 43], or providing guarantees on relevant risks leveraging p-values or risk estimates obtained from concentration inequalities [3, 7, 4]. Such works differ from our two-step conformal approach in several ways, such as (i) considering different vision tasks and using different data modalities, (ii) integrating CP into complex modelling pipelines that cease to be post-hoc and model-agnostic, or (iii) leveraging non-CP methods.
CP intervals for bounding boxes have been previously considered by Andéol et al. [14, 1], whom we consider closest prior work. A key limitation in their approach is that only a single object class is considered, thus providing the most trivial form of our guarantee in Equation 3. In particular, the misclassification risk we address in section 5 does not arise for the single-class setting, where the class label is known a priori and thus $\hat{l}_{n+1}=l_{n+1}$ trivially holds. This strongly limits the practicality of their approach for real-world settings where multiple classes interact, such as in autonomous driving. Thus, our provided guarantees are theoretically stronger and more practically relevant.
In addition, we introduce several methodological improvements with good results (see subsection 6.2), such as (i) novel ensemble and quantile CP adaptations that move beyond their proposed scoring approaches, (ii) more informative two-sided intervals, as opposed to their simpler one-sided intervals, and (iii) a multiple testing correction that effectively exploits correlation structure between box coordinates, as opposed to a more naive Bonferroni [14] or $\max$-correction [1].

### v2

@@ Abstract
Quantifying a model’s predictive uncertainty is essential for safety-critical applications such as autonomous driving. We consider quantifying such uncertainty for multi-object detection. In particular, we leverage conformal prediction to obtain uncertainty intervals with guaranteed coverage for object bounding boxes. One challenge in doing so is that bounding box predictions are conditioned on the object’s class label. Thus, we develop a novel two-step conformal approach that propagates uncertainty in predicted class labels into the uncertainty intervals of bounding boxes. This broadens the validity of our conformal coverage guarantees to include incorrectly classified objects, thus offering more actionable safety assurances. Moreover, we investigate novel ensemble and quantile regression formulations to ensure the bounding box intervals are adaptive to object size, leading to a more balanced coverage. Validating our two-step approach on real-world datasets for 2D bounding box localization, we find that desired coverage levels are satisfied with practically tight predictive uncertainty intervals.
Keywords:
Object Detection Conformal Prediction Uncertainty
@@ 1 Introduction
Safety-critical applications in domains such as autonomous transportation [37, 69] and mobile robotics [36] benefit greatly from accurate estimates of the model’s predictive uncertainty. Yet one obstacle to principled uncertainty quantification (UQ) for computer vision is the pervasive use of deep neural networks, which are often unamenable to traditional techniques for UQ. The framework of Conformal Prediction (CP) [65, 54, 2] enables a form of distribution-free UQ that is agnostic to the predictive model’s structure, rendering it well-suited for such ‘black-box’ models.
In this work, we propose a CP framework designed to quantify predictive uncertainties in multi-object detection tasks with multiple classes (see Fig. 2). CP allows us to produce computationally cheap, post-hoc distribution-free prediction intervals, which come equipped with a coverage guarantee for the true bounding boxes of previously unseen objects (of known classes). Specifically, we provide users with the following safety assurance: “The conformal prediction interval covers the object’s true bounding box with probability $(1-\alpha)$ for any known object class”, where $\alpha$ is an acceptable margin of error. Such a guarantee can, e.g., in the context of autonomous driving, help certify collision avoidance by steering clear of the outer interval bounds, or in the case of robot picking, enforce cautious handling by demarcating a reliable grasping zone via the inner bounds. We provide visual examples of our obtained intervals in Fig. 1 and § 0.C.5.
@@
Employing strategies based on ensembling and quantile regression, we ensure that the obtained intervals are adaptive to object size: they may grow or shrink in individual dimensions to account for object variability and prediction difficulty. A challenge to the desired assurance is that constructed intervals rely on the model’s predicted class labels, which may be erroneous. We thus introduce an additional conformal step over the class labels, shielding against misclassification and ensuring that downstream coverage is satisfied. That is, our two-step conformal pipeline remains theoretically and empirically valid regardless of the underlying object detector’s predictive performance for either class labels or box coordinates – the incurred costs are solely reflected in the obtained prediction interval sizes. In the experiments, we apply our methodology to multiple classes on several real-world 2D object detection datasets. We obtain bounding box prediction intervals that adhere to the desired guarantee, and are both adaptive and practically useful for downstream decision-making.
Figure 2: A diagram of our proposed two-step conformal approach. We compute conformal quantiles for both class labels and box coordinates on calibration data following the CP framework. These are used on the predictions of a ‘black-box’ object detector for a new test sample to (1) form a conformal label set with guarantee (✓) which informs our box quantile choice, and (2) form a conformal prediction interval for the bounding box with guarantee (✓), providing a reliable predictive uncertainty estimate.
To summarize, our core contribution is an end-to-end framework for safe bounding box uncertainties which is post-hoc, efficient, and generalizable. In that process, we introduce several original concepts such as (i) ensemble and quantile CP adaptations for object detection, (ii) leveraging strong class-conditional guarantees for multi-class settings, and (iii) proposing a sequential two-step approach that propagates classification uncertainties forward.
@@ 2 Background
We begin by providing background on conformal prediction and the desired coverage guarantees, and then relate our object detection setting to it.
@@ 2.1 Conformal prediction
We consider the most common setting of split CP [43], where we perform a single split to obtain hold-out calibration data $\mathcal{D}_{cal}=\{(X_{i},Y_{i})\}_{i=1}^{n}\sim P_{XY}$, as opposed to alternative partitioning schemes [64, 6]. If the general conformal procedure outlined in Algorithm 1 (deferred to § 0.A.1) is followed, a coverage guarantee for an unseen test sample $(X_{n+1},Y_{n+1})\sim P_{XY}$ is provided in terms of a prediction set $\hat{C}(X_{n+1})$, where a finite-sample, distribution-free guarantee is given over the event of $\hat{C}(X_{n+1})$ containing $Y_{n+1}$.
That is, assuming the samples $\mathcal{D}_{cal}\cup\{(X_{n+1},Y_{n+1})\}$ are exchangeable – a relaxation of the i.i.d. assumption – we obtain a probabilistic guarantee that
@@
for some tolerated miscoverage rate $\alpha\in(0,1)$ [54]. The provided guarantee is marginally valid, since it holds on average across any sample $(X_{n+1},Y_{n+1})$ and set $\mathcal{D}_{cal}$ drawn from some fixed distribution $P_{XY}$ over $\mathcal{X}\times\mathcal{Y}$. This is in contrast to the ideal scenario of conditionally valid coverage per input $X_{n+1}$, which has been shown to be impossible to achieve in a distribution-free manner [63, 19]. However, recent work on in-between notions of conditionality such as group- [49, 23] and feature-conditional [53] strive towards more granular guarantees.
@@
where $\mathcal{Y}=\{1,\dots,K\}$ are distinct class labels. The class-conditional guarantee in Eq. 2 is stronger and implies Eq. 1, in that we aim to control the miscoverage level for samples within each class. It also permits setting individual miscoverage levels $\{\alpha_{y}\}_{y\in\mathcal{Y}}$ per class if desired, and is robust to imbalances in class proportions [47, 62]. Such a class-conditional guarantee is precisely what we aim to provide.
Classification and regression. Applying CP to a classification task yields conformal label prediction sets $\hat{C}_{L}(X_{n+1})\subseteq\{1,\dots,K\}$ as finite subsets of the $K$ class labels, at a target miscoverage level $\alpha_{L}$. For regression, the sets $\hat{C}_{B}(X_{n+1})\subseteq\mathbb{R}$ take the form of prediction intervals (PIs) on the target domain, at a target miscoverage level $\alpha_{B}$. Naturally, we have both $(\alpha_{L},\alpha_{B})\in(0,1)^{2}$.
@@ 2.2 Object detection
We next formalize our multi-object detection setting. Consider an input image $X\in\mathbb{R}^{H\times W\times D}$, where $H$, $W$ and $D$ correspond to image height, width and depth. For each image in $\mathcal{D}_{cal}$ we also receive a set of tuples $(c^{1},c^{2},c^{3},c^{4},l)$, where $(c^{1},c^{2},c^{3},c^{4})\in\mathbb{R}^{4}$ are the coordinates indicating an object’s bounding box location in the image, and $l\in\{1,\dots,K\}$ represents the object’s class label.
Each tuple parameterizes an object, with a total of $O(X)$ true objects located in the image. For image $X$ we thus have objects $\{(c^{1},c^{2},c^{3},c^{4},l)_{j}\}_{j=1}^{O(X)}$. Note that the model predicts $\hat{O}(X)$ objects, and it is possible that $O(X)\neq\hat{O}(X)$. We model every object as an individual sample for our CP procedures, i.e., the same input image $X$ can produce multiple calibration samples of shape $(X,(c^{1},c^{2},c^{3},c^{4},l)_{j})$, where $j=1,\dots,O(X)$ denote the contained objects.
Object detection model. For our object detector $\hat{f}$, we define two separate output heads. The probabilistic classification head is defined as the map $\hat{f}_{L}:X\mapsto(\hat{\pi}_{1},\dots,\hat{\pi}_{K})$, where $\hat{\pi}_{y}$ is the model’s estimate of the true class probability $\pi_{y}$ of some object in image $X$ belonging to class $y$. The object’s class label is then $l=\arg\max_{y\in\{1,\dots,K\}}\,\hat{\pi}_{y}$. The bounding box regression head, denoted as $\hat{f}_{B}:X\mapsto(\hat{c}^{1},\hat{c}^{2},\hat{c}^{3},\hat{c}^{4})$, maps to an object’s real-valued bounding box coordinates.
@@ 2.3 Conformal prediction for object detection
Given our multi-object detection setting, we consider a class-conditional CP approach to be particularly meaningful. It is sensible to only leverage information on detected objects of the same class, e.g., class ‘car’, to construct PIs for new objects of that class. In contrast, a general marginal approach will unintuitively also employ information from unrelated classes, such as ‘person’ or ‘bicycle’.
We apply CP to the bounding boxes on a per-coordinate basis, previously denoted $(c^{1},c^{2},c^{3},c^{4})$. However, from now on let us consider the generalization to an arbitrary amount of coordinates $c^{k},\,k=1,\dots,m$  [^1] This easily permits extending our approach to higher-dimensional object parameterizations such as 3D bounding boxes.. If we consider the class label $l$ within each group of objects belonging to a common class as fixed (since the same label is shared), the response of an individual sample $(X_{i},Y_{i})$ can be interpreted as a realization of the $m$ coordinates only, i.e., we define $Y_{i}:=(c_{i}^{1},\dots,c_{i}^{m})\in\mathbb{R}^{m}$. The desired guarantee in Eq. 2 is re-interpreted as
$\mathbb{P}\left(\bigcap_{k=1}^{m}\left(c_{n+1}^{k}\in\hat{C}_{B}^{k}(X_{n+1})\right)\,|\,l_{n+1}=y\right)\geq 1-\alpha_{B}\quad\forall y\in\mathcal{Y},$ (3)
where components are indexed accordingly per specific coordinate dimension. For example, $\hat{C}_{B}^{k}(X_{n+1})$ is the $k$-th coordinate’s prediction interval of an object of class $y$ (its class label $l_{n+1}$ matches $y$) located in image $X_{n+1}$. Applying CP per coordinate gives rise to multiple testing issues, which we address in § 4.1.
Practical limitations. Naively applying class-conditional CP to the box coordinates necessitates a correct class label prediction in order to satisfy validity. That is, for Eq. 3 to hold, a valid PI construction requires $\hat{l}_{n+1}=l_{n+1}$ for any considered class $y\in\mathcal{Y}$. We alleviate this practically limiting dependence on the model’s classification ability using a conformal set-based classifier in § 5 (see also Fig. 2). However, we of course still rely on the model’s general detection abilities: the provided guarantees only hold for true objects that are actually detected (true positives) and do not account for undetected objects (false negatives), as also noted by Andéol et al. [15, 1]. Finally, the assumption on data exchangeability underlying CP requires $P_{XY}$ to remain fixed, albeit recent works have explored CP under settings of mild or known distribution shifts [2, 18].
@@ 3 Related work
Many existing approaches for uncertainty in bounding box regression leverage standard UQ techniques such as Bayesian inference [13, 25, 71], loss attenuation [22, 29, 30], or practical approximations like Monte Carlo Dropout [45, 21, 39, 74] and Deep Ensembles [16, 68, 40]. These can require substantial modifications to the model architecture or training procedure, and do not provide a guarantee or statement of assurance about provided estimation quality. See Feng et al. [17] for a recent survey. A complementary line of work investigates the calibration of object detectors [46, 27, 42], which can benefit our approach by improving the underlying ‘black-box’ probabilistic model.
Conformal approaches have recently gained traction for computer vision and related tasks, with applications such as image classification [51, 5], geometric pose estimation [73], or tracking and trajectory planning [58, 34, 35, 41]. Yet, the domain remains comparatively unexplored given current surveys [2, 18]. Specific attempts at principled UQ for bounding boxes include using the Probably Approximatly Correct (PAC) framework to produce guarantees by composition of PAC sets at multiple modelling stages [32, 44], or leveraging p-values and risk estimates obtained from concentration inequalities for related risk control [3, 7, 4]. Such works differ from our two-step conformal approach in several ways, such as (i) considering different vision tasks and using different data modalities, (ii) integrating CP into complex modelling pipelines that cease to be post-hoc and model-agnostic, or (iii) employing methods not based on conformal prediction.
Closest prior work. Conformal PIs for bounding boxes have been previously considered by Andéol et al. [15, 1]. However, a crucial limitation in their approach is that bounding box uncertainty is considered and evaluated for a single class and for correctly classified objects only. Thus only the simplest form of our guarantee in Eq. 3 is provided, since the class label is known a priori and therefore $\hat{l}_{n+1}=l_{n+1}$ trivially holds. This means that prevalent uncertainty in the class label predictions (which we address in § 5) is entirely ignored, making their approach unsuitable for settings with multiple interacting classes, such as autonomous driving. We also introduce several methodological improvements, such as (i) novel ensemble and quantile scoring functions for the bounding box setting, (ii) more informative two-sided intervals, and (iii) a multiple testing correction that exploits correlation structure between box coordinates, as opposed to more naive Bonferroni [15] or $\max$-corrections [1].